Acids have been made famous by their use in spy movies. James Bond often carries an acid-filled gold pen. When Bond is captured and imprisoned he squirts some acid out of his pen and onto the iron bars of the cell. The acid quickly dissolves the metal allowing Bond to escape. What minimum volume of 12.00M hydrochloric acid would be required to completely dissolve a 500.00g iron bar? What aspects of this situation are not feasible?
Fe(s) + 2HCl(aq) ⇨ FeCl2(aq) + H2(g)
1
Expert's answer
2020-07-07T14:20:48-0400

 Let minimum volume of 12.00 M hydrochloric acid would be required to completely dissolve a 500.00 g iron (Fe) . is " V" liter .


So , Moles of HCl = ( V * 12 ) mole = 12 V mole ..................1.


Moles of Fe = Mass of Fe / Molar Mass of Fe = 500 / 56 = 8.92857 . mole .


Given Reaction is


Fe(s) + 2HCl(aq) ⇨ FeCl2(aq) + H2(g) ..............


1 mole of Fe Completely React With 2 mole of HCl.


8.92857 mole of Fe Completely React With ( 2*8.92857 ) = 17.85714 mole of HCl..............


So moles of HCl Reacted = 17.85714 mol. ..............................2..



From 1 and 2 Both are Represent Moles of HCl in This Reaction So Both Should Be Equal .


12 V mole = 17.85714 mole


V =17.85714 / 12 . = 1. 488 liter .